OpenStudy (dls):

What is tan^2 ( arcsin 1) = ?

OpenStudy (dls):

Infinity or undefined?

OpenStudy (dls):

@dan815 @genius12 @agent0smith

OpenStudy (anonymous):

\[\bf \arcsin(1)=\frac{ \pi }{ 2 }\]

OpenStudy (dls):

so undefined whole square shud be undef again?

OpenStudy (anonymous):

\[\bf (\tan(\pi/2))^2=??\]

OpenStudy (dls):

undef?

OpenStudy (anonymous):

correct lol or \(\bf +\infty\)

OpenStudy (dls):

just wanted to confirm,thanks! also infinity whole square is infinity?

OpenStudy (anonymous):

ya lol

OpenStudy (anonymous):

tan^2(x) still has a vertical asymptote at x = pi/2 just like tan(x) hence as x tends to pi/2, the function tends to positive infinity.

OpenStudy (dls):

thanks :P

OpenStudy (dan815):